OpenStudy (anonymous):

Mr. Cruz cut three pieces of wood for a project. The first piece of lumber is the longest. The second piece is 7/8 times as long as the first piece. The third piece is 12 5/16 inches shorter than the second piece. The third piece is 24 7/16 inches long. a. How long is the first piece of wood? b. How much wood did Mr. Cruz use to make all 3 pieces? c. Mr. Cruz has 32 feet of lumber left. He wants to cut 4 more of each of the pieces described above. Does he have enough lumber? Justify the answer.

OpenStudy (anonymous):

x = first piece y = 2nd piece z = 3rd piece x is the longest y = 7/8x z = y - 12 5/16 which = 7/8x - 12 5/16 z = 24 7/16 7/8x - 12 5/16 = 24 7/16 7/8x - 197/16 = 391/16 7/8x = 391/16 + 197/16 7/8x = 588/16 x = 588/16 * 8/7 x = 4704/112 x = 42 y = 7/8x y = 7/8(42) y = 294/8 y = 36 3/4 x = 42 y = 36 3/4 z = 24 7/16 any questions ?

OpenStudy (anonymous):

a) first piece of wood is 42 b) all three pieces (42 + 36 3/4 + 24 7/16) 42 + 36 + 24 = 102 3/4 + 7/16 = 12/16 + 7/16 = 19/16 = 1 3/26 102 + 1 3/16 = 103 3/16 <-- total c) he has 32 feet of lumber 4(103 3/16) = 4 * 1651/16 = 6604/16 = 412 3/4 412.75/12 = 34.40 ft. He does not have enough